# FELLE: Autoregressive Speech Synthesis with Token-Wise Coarse-to-Fine Flow Matching

### v1

@@ Abstract
To advance continuous-valued token modeling and temporal-coherence enforcement, we propose FELLE, an autoregressive model that integrates language modeling with token-wise flow matching. By leveraging the autoregressive nature of language models and the generative efficacy of flow matching, FELLE effectively predicts continuous-valued tokens (mel-spectrograms). For each continuous-valued token, FELLE modifies the general prior distribution in flow matching by incorporating information from the previous step, improving coherence and stability. Furthermore, to enhance synthesis quality, FELLE introduces a coarse-to-fine flow-matching mechanism, generating continuous-valued tokens hierarchically, conditioned on the language model’s output. Experimental results demonstrate the potential of incorporating flow-matching techniques in autoregressive mel-spectrogram modeling, leading to significant improvements in TTS generation quality, as shown in https://aka.ms/felle.
@@
Figure 1: Overview of FELLE, an autoregressive mel-spectrograms model that generates personalized speech from text and acoustic prompts. At each timestep, the framework relies on the previous mel-spectrogram distribution as a prior, conditioned on the output of the language model, applying a coarse-to-fine flow-matching module to produce refined spectral features.
1 Introduction
@@
The remarkable success of large language models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Team et al., 2024) has prompted a paradigm shift in speech synthesis, redefining it as a language modeling task. This shift has driven notable progress in zero-shot speech synthesis (Chen et al., 2025, 2024a). Consistent with the standard LLM training methodology, researchers have naturally adopted discrete-valued tokens as the foundational modeling units. However, unlike textual data, which is inherently discrete, speech signals require complex quantization techniques to transform continuous waveforms into discrete-valued tokens. These essential quantization processes impose fundamental constraints compared to continuous representations, particularly in terms of fidelity preservation and training complexity (Puvvada et al., 2024; Meng et al., 2024). Consequently, discrete token-based text-to-speech (TTS) systems often face challenges such as intricate modeling workflows and reduced output quality. In response to these limitations, recent research has increasingly explored autoregressive (AR) modeling frameworks that leverage continuous representations (Meng et al., 2024; Turetzky et al., 2024; Zhu et al., 2024), showing notable improvements in model performance and simplifying training processes.
@@
2 Related Work and Background
2.1 Related Work
@@
Zero-shot text-to-speech approaches are commonly categorized into autoregressive and non-autoregressive paradigms based on their output generation mechanisms. Autoregressive systems typically rely on language model architectures (Chen et al., 2025; Kharitonov et al., 2023; Yang et al., 2024), whereas non-autoregressive implementations commonly employ diffusion models and analogous methodologies (Ju et al., 2024; Chen et al., 2024b). The subsequent discussion concentrates on research efforts investigating diverse representations under the framework of autoregressive language modeling architectures.
Discrete-Valued Token-Based TTS
@@
TTS systems based on discrete representations utilize tokenized acoustic units derived from unsupervised or semi-supervised learning frameworks. These discrete tokens serve as compact and efficient representations of speech, capturing phonetic and prosodic attributes while reducing redundancy in data storage and computation. VALL-E (Chen et al., 2025) is a neural codec language model for text-to-speech synthesis that firstly redefines TTS as a conditional language modeling task, enabling high-quality, personalized speech generation from just a 3-second acoustic prompt, significantly advancing naturalness and speaker similarity. Recent studies further enhance VALL-E’s capabilities across multilingual generalization (Zhang et al., 2023), decoding efficiency (Chen et al., 2024a), and robustness (Song et al., 2024; Xin et al., 2024; Han et al., 2024), collectively advancing zero-shot speech synthesis in scalability, quality, and linguistic flexibility. In contrast to the unified language modeling approach of VALL-E and its variants, CosyVoice (Du et al., 2024) leverages an LLM for text-to-token conversion followed by a conditional flow-matching model for token-to-spectrogram synthesis, enhancing zero-shot voice cloning through end-to-end supervised speech token learning.
Continuous-Valued Token-Based TTS
@@
Recent advances in continuous representation-based TTS systems eliminate the need for cumbersome codec training while achieving promising performance. Notably, MELLE (Meng et al., 2024) proposes a single-pass language model architecture leveraging rich continuous acoustic representations, enabling precise control over prosodic features including pitch, rhythm, and timbre for high-fidelity speech synthesis. In contrast, SALAD (Turetzky et al., 2024) is a zero-shot text-to-speech system that employs a per-token latent diffusion model on continuous representations, enabling variable-length audio generation through semantic tokens for contextual guidance and stopping control. While this method achieves superior intelligibility scores, it may face challenges related to time costs. Alternatively, KALL-E (Zhu et al., 2024) adopts an autoregressive approach with WaveVAE to directly model speech distributions, bypassing both VAE and diffusion paradigms, demonstrating enhanced naturalness and speaker similarity through probabilistic waveform prediction.
2.2 Background
@@
Flow matching (Lipman et al., 2022) is a technique for learning a transformation that maps a prior distribution $p_{0}$ to a target distribution $q(x)$. The core idea of flow matching is to define a flow $\phi_{t}(x)$ that evolves over time, transforming the prior distribution $p_{0}$ into the target distribution $q(x)$. This flow $\phi_{t}(x)$ is governed by a vector field $v_{t}(x)$ and satisfies the following ordinary differential equation (ODE):
$\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\quad\phi_{0}(x)=x.$ (1)
Here, $\phi_{0}(x)=x$ indicates that at time $t=0$, the flow $\phi_{t}(x)$ is an identity mapping.
While flow matching provides a principled framework for learning such transformations, it can be computationally expensive due to the difficulty of directly accessing the true vector field $u_{t}(x)$ and the target distribution $q(x)$. To address this, Conditional Flow Matching (CFM) is introduced. In CFM, the flow and the vector field are conditioned on the data $x_{1}$, making the optimization process more efficient. The objective of CFM is to minimize the discrepancy between the conditional true vector field $u_{t}$ and the learned conditional vector field $v_{t}(x;\theta)$. This discrepancy is measured by the following loss function:
$L_{\text{CFM}}=\mathbb{E}_{t,x_{1},x}\left\|u_{t}-v_{t}(x;\theta)\right\|^{2},$ (2)
where time $t$ is uniformly sampled from $\mathcal{U}[0,1]$, data points $x_{1}$ are drawn from the target distribution $q(x_{1})$, samples $x$ are generated through the conditional probability path $p_{t}(x|x_{1})$, and the conditional vector field $u_{t}\equiv u_{t}(x|x_{1})$.
3 FELLE
3.1 Problem Formulation: Token-wise Flow Matching for AR Model
@@
$p(\bm{X}\!\mid\!\bm{y})\!$ $=\prod_{i=0}^{L-1}p(\bm{x}^{i}\mid\bm{x}^{<i},\bm{y},\bm{\widehat{x}})$ (3)
$=\!\prod_{i=0}^{L-1}p_{\theta_{\text{FM}}}(\bm{x}^{i}\mid\bm{z}^{i}),\bm{z}^{i}\!=\!f_{\theta_{\text{LM}}}(\bm{x}^{<i},\bm{y},\bm{\widehat{x}}).$
@@
3.2 FELLE Architecture
The proposed framework combines an autoregressive language model with a flow-matching mechanism, which facilitates the progressive generation of high-fidelity speech. As shown in Figure 1, the autoregressive model $f_{\theta_{\text{LM}}}$ extracts features from the text prompt $\bm{y}$ and speech prompt $\bm{\widehat{x}}$, generating latent representations $\bm{z}^{i}$ (where $i$ denotes the generation step) that serve as conditional inputs for the flow-matching mechanism. The flow-matching mechanism applies a coarse-to-fine strategy to generate high-quality mel-spectrogram frames $\bm{x}^{i}$. The main components of the approach are described in detail below.
Figure 2: The coarse-to-fine flow-matching module of FELLE. (a) The training process along with the detailed data flow within the coarse-to-fine module. The gray dashed lines merely indicate the relationships between components in the model structure and are not activated during training. (b) The inference process.
3.2.1 Autoregressive Language Model
The language model, designed as a unidirectional Transformer decoder, generates acoustic features autoregressively by utilizing both text sequences and mel-spectrogram prompts. In the initial step, the text tokens are embedded, while a pre-net maps the mel-spectrogram into the dimensional space of the LM. By processing the combined text $\bm{y}$, speech prompt $\bm{\widehat{x}}$, and acoustic embeddings $\bm{x}^{<i}$, the language model $f_{\theta_{\text{LM}}}$ processes multi-head attention and feed-forward layers to capture the intricate relationship between linguistic and acoustic information. The output at each time step subsequently serves as a conditioning input for the coarse-to-fine flow-matching module to synthesize the next-frame acoustic features.
3.2.2 Coarse-to-Fine Flow Matching
@@
In the fine stage, the model refines this approximation by recovering details $x^{i,f}$, represented as the residual between the original frame $x^{i}$ and the upsampled coarse component $\text{{Upsample}}(x^{i,c})$. A secondary flow-matching model predicts the vector field $v_{t}^{f}(x^{i,f},z^{i},x^{i,c};\theta_{\text{FM}}^{f})$, governing this process by leveraging both the features $z^{i}$ and the coarse component (with ground-truth coarse features $x^{i,c}$ during training and predicted values $\tilde{x}^{i,c}$ during inference) as conditional inputs. This hierarchical conditioning allows the fine model to focus on local details while preserving global coherence from the coarse stage.
@@
$\mathcal{L}_{\text{C2F-FM}}=\underbrace{\mathbb{E}_{t,x_{1}^{i,c},x^{i,c}}\left\|u_{t}^{c}-v_{t}^{c}(x^{i,c},z^{i};\theta_{\text{FM}}^{c})\right\|^{2}}_{\text{Coarse Stage}}+\underbrace{\mathbb{E}_{t,x_{1}^{i,f},x^{i,f}}\left\|u_{t}^{f}-v_{t}^{f}(x^{i,f},z^{i},x_{1}^{i,c};\theta_{\text{FM}}^{f})\right\|^{2}}_{\text{Fine Stage}},$ (5)
@@
Classifier-free guidance (CFG) is a powerful technique to enhance the quality and controllability of generated outputs in flow matching and diffusion models (Ho and Salimans, 2022; Nichol and Dhariwal, 2021). In FELLE, we implement CFG through joint training of coarse and fine flow matching models using both conditional and unconditional objectives. During training, we randomly mask the speech prompt with probability $p_{\text{drop}}$ for unconditional learning, which enables each model to learn dual vector fields.
At inference, guided vector fields are computed through linear blending:
$\hat{v}_{t}^{\ast}(x^{\ast};\cdot)$ $=wv_{t}^{\ast}(x^{\ast},c;\theta_{\text{FM}}^{\ast})+(1-w)v_{t}^{\ast}(x^{\ast},\bar{c};\theta_{\text{FM}}^{\ast}),$ (6)
@@
3.3 Training Objective
In FELLE, we integrate the condition loss $\mathcal{L}_{\text{cond}}$ in addition to coarse-to-fine loss $\mathcal{L}_{\text{C2F-FM}}$. $\mathcal{L}_{\text{cond}}$ is a hybrid loss function that combines L1 and L2 norms, defined as $\mathcal{L}_{\text{cond}}=\|z_{i}-x_{i}\|_{1}+\|z_{i}-x_{i}\|_{2}^{2}$, for step $i$ to regularize the conditional input for flow matching. Additionally, we introduce a stop prediction module to the autoregressive language model. This module, during each step of generation, transforms the hidden state output by the language model into the probability of a stop signal through a linear layer and calculates the Binary Cross-Entropy loss $\mathcal{L}_{\text{stop}}$ for training. The model can automatically determine when to stop during the generation process without the need to preset length rules. The overall training objective is:
$\mathcal{L}=\mathcal{L}_{\text{C2F-FM}}+\lambda\mathcal{L}_{\text{cond}}+\alpha\mathcal{L}_{\text{stop}}$ (7)
where $\lambda$ and $\alpha$ control the respective contributions of $\mathcal{L}_{\text{cond}}$ and $\mathcal{L}_{\text{stop}}$.
3.4 Inference
As illustrated in Figure 2(b), the inference process employs an autoregressive language model that progressively generates hidden representations based on textual and speaker prompts. At each step $i$, the computed latent state $z_{i}$ serves two key purposes. First, it provides conditional guidance for the coarse flow-matching module, facilitating the gradual transformation from the previous mel-spectrogram approximation $\tilde{x}^{i-1,c}$ to the current coarse structural estimate $\tilde{x}^{i,c}$. Following this coarse estimation phase, the integrated information of $\tilde{x}^{i,c}$ and $z_{i}$ drives the fine flow-matching module to produce the fined mel-spectrogram frame $\tilde{x}^{i,f}$. The final output frame $\tilde{x}^{i}$ emerges through the integration of these complementary coarse and refined predictions. Secondly, the latent state $z_{i}$ processed by the stop prediction module to compute the stop probability, which is compared against a predefined threshold to decide whether to terminate the process. The iterative generation continues until the stop criterion is satisfied, after which a neural vocoder converts the accumulated mel-spectrogram into the final speech waveform.
4 Experimental Setup
4.1 Dataset

### v2

@@ Abstract.
To advance continuous token modeling and temporal-coherence enforcement, we propose FELLE, an autoregressive model that integrates language modeling with token-wise flow matching. By leveraging the autoregressive nature of language models and the generative efficacy of flow matching, FELLE effectively predicts continuous-valued tokens (mel-spectrograms). For each continuous-valued token, FELLE modifies the general prior distribution in flow matching by incorporating information from the previous step, improving coherence and stability. Furthermore, to enhance synthesis quality, FELLE introduces a coarse-to-fine flow-matching mechanism, generating continuous-valued tokens hierarchically, conditioned on the language model’s output. Experimental results demonstrate the potential of incorporating flow-matching techniques in autoregressive mel-spectrogram modeling, leading to significant improvements in TTS generation quality, as shown in https://aka.ms/felle.
Zero-shot Text-to-Speech, Autoregressive Modeling, Continuous-valued Token Modeling, Coarse-to-Fine Generation, Flow Matching
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755494isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologiesccs: Computing methodologies Natural language generation
@@
Figure 1. FELLE is an autoregressive mel-spectrogram model that generates personalized speech from text and acoustic prompts. It uses the previous mel-spectrogram as a prior and refines features with a coarse-to-fine flow-matching module guided by the language model.
[…]
Diagram showing the architecture of FELLE, an autoregressive model that generates personalized speech from text and acoustic prompts. The model predicts mel-spectrograms by conditioning on both previous outputs and language model features, using a coarse-to-fine flow-matching module for refinement.
1. Introduction
@@
The remarkable success of large language models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Team et al., 2024) has prompted a paradigm shift in speech synthesis, redefining it as a language modeling task. This shift has driven notable progress in zero-shot speech synthesis (Chen et al., 2025, 2024a). Consistent with the standard LLM training methodology, researchers have naturally adopted discrete-valued tokens as the foundational modeling units. However, unlike textual data, which is inherently discrete, speech signals require complex quantization techniques to transform continuous waveforms into discrete-valued tokens. These essential quantization processes impose fundamental constraints compared to continuous representations, particularly in terms of fidelity preservation and training complexity (Puvvada et al., 2024; Meng et al., 2024). Consequently, discrete token-based text-to-speech (TTS) systems often face challenges such as intricate modeling workflows and reduced output quality. In response to these limitations, recent research has increasingly explored autoregressive (AR) modeling frameworks that leverage continuous representations (Meng et al., 2024; Turetzky et al., 2024; Zhu et al., 2024; Ma et al., 2025), showing notable improvements in model performance and simplifying training processes.
@@
2. Related Work
@@
Zero-shot TTS are commonly categorized into autoregressive and non-autoregressive paradigms based on their output generation mechanisms. Autoregressive systems typically rely on language model architectures (Chen et al., 2025; Kharitonov et al., 2023; Yang et al., 2024), whereas non-autoregressive implementations commonly employ diffusion models and analogous methodologies (Ju et al., 2024; Chen et al., 2024b; Wang et al., 2025a). The subsequent discussion focuses on research efforts investigating diverse representations under the framework of autoregressive language modeling architectures.
2.1. Discrete-Valued Token-Based TTS
@@
TTS systems based on discrete representations (Chen et al., 2025; Łajszczak et al., 2024; Song et al., 2024; Du et al., 2024a, b) utilize tokenized acoustic units derived from unsupervised or semi-supervised learning frameworks. These discrete tokens serve as compact representations of speech, capturing phonetic and prosodic attributes while reducing redundancy in data storage and computation. VALL-E (Chen et al., 2025) is a neural codec language model for text-to-speech synthesis that firstly redefines TTS as a conditional language modeling task, enabling high-quality, personalized speech generation from just a 3-second acoustic prompt, significantly advancing naturalness and speaker similarity. Recent studies further enhance VALL-E’s capabilities across multilingual generalization (Zhang et al., 2023), decoding efficiency (Chen et al., 2024a), and robustness (Song et al., 2024; Xin et al., 2024; Han et al., 2024), collectively advancing zero-shot TTS in scalability, quality, and linguistic flexibility. In contrast to the unified language modeling approach of VALL-E and its variants, CosyVoice (Du et al., 2024a) leverages an LLM for text-to-token conversion followed by a conditional flow-matching model for token-to-spectrogram synthesis, enhancing zero-shot voice cloning through end-to-end supervised speech token learning.
2.2. Continuous-Valued Token-Based TTS
@@
Recent advances in continuous representation-based TTS systems eliminate the need for cumbersome codec training while achieving promising performance. Notably, MELLE (Meng et al., 2024) proposes a single-pass language model architecture leveraging rich continuous acoustic representations, enabling precise control over prosodic features including pitch, rhythm, and timbre for high-fidelity speech synthesis. SALAD (Turetzky et al., 2024) is a zero-shot text-to-speech system that employs a per-token latent diffusion model on continuous representations, enabling variable-length audio generation through semantic tokens for contextual guidance and stopping control. While this method achieves superior intelligibility scores, it may face challenges related to time costs. Alternatively, KALL-E (Zhu et al., 2024) adopts an autoregressive approach with WaveVAE to directly model speech distributions, bypassing both VAE and diffusion paradigms, demonstrating enhanced performance through probabilistic waveform prediction.
3. Preliminary
3.1. Background
@@
Flow matching (Lipman et al., 2022) is a technique for learning a transformation that maps a prior distribution $p_{0}$ to a target distribution $q(x)$. The core idea of flow matching is to define a flow $\phi_{t}(x)$ that evolves over time, transforming the prior distribution $p_{0}$ into the target distribution $q(x)$. This flow $\phi_{t}(x)$ is governed by a vector field $v_{t}(x)$ and satisfies the following ordinary differential equation: $\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\quad\phi_{0}(x)=x.$ Here, $\phi_{0}(x)=x$ indicates that at time $t=0$, the flow $\phi_{t}(x)$ is an identity mapping.
While flow matching provides a principled framework for learning such transformations, it can be computationally expensive due to the difficulty of directly accessing the true vector field $u_{t}(x)$ and the target distribution $q(x)$. To address this, Conditional Flow Matching (CFM) is introduced. In CFM, the flow and the vector field are conditioned on the data $x_{1}$, making the optimization process more efficient. The objective of CFM is to minimize the discrepancy between the conditional true vector field $u_{t}$ and the learned conditional vector field $v_{t}(x;\theta)$. This discrepancy is measured by the following loss function: $L_{\text{CFM}}=\mathbb{E}_{t,x_{1},x}\left\|u_{t}-v_{t}(x;\theta)\right\|^{2},$ where time $t$ is uniformly sampled from $\mathcal{U}[0,1]$, data points $x_{1}$ are drawn from the target distribution $q(x_{1})$, samples $x$ are generated through the conditional probability path $p_{t}(x|x_{1})$, and the conditional vector field $u_{t}\equiv u_{t}(x|x_{1})$.
3.2. Problem Formulation
@@
[…]
@@
4. FELLE Architecture
The proposed framework combines an autoregressive language model with a flow-matching mechanism, which facilitates the progressive generation of high-fidelity speech. As shown in Figure 1, the autoregressive model $f_{\theta_{\text{LM}}}$ extracts features from the text prompt $\bm{y}$ and speech prompt $\bm{\widehat{x}}$, generating latent representations $\bm{z}^{i}$ (where $i$ denotes the generation step) that serve as conditional inputs for the flow-matching mechanism. The flow-matching mechanism applies a coarse-to-fine strategy to generate high-quality mel-spectrogram frames $\bm{x}^{i}$. The main components are described in detail below.
Figure 2. The coarse-to-fine flow-matching module of FELLE. (a) The training process along with the detailed data flow within the coarse-to-fine module. The gray dashed lines merely indicate the relationships between components in the model structure and are not activated during training. (b) The inference process.
[…]
Two-part diagram of the coarse-to-fine flow-matching module in FELLE. Part (a) shows the training process, illustrating the internal data flow through the coarse and fine stages. Gray dashed lines represent inactive connections used only for model structure reference. Part (b) depicts the inference process with active data paths.
4.1. Autoregressive Language Model
The language model, designed as a Transformer decoder, generates acoustic features autoregressively by utilizing both text sequences and mel-spectrogram prompts. In the initial step, the text tokens are embedded, while a pre-net maps the mel-spectrogram into the dimensional space of the LM. By processing the combined text $\bm{y}$, speech prompt $\bm{\widehat{x}}$, and acoustic embeddings $\bm{x}^{<i}$, the language model $f_{\theta_{\text{LM}}}$ processes multi-head attention and feed-forward layers to capture the intricate relationship between linguistic and acoustic information. The output at each time step subsequently serves as a conditioning input for the coarse-to-fine flow-matching module to synthesize the next-frame acoustic features.
4.2. Coarse-to-Fine Flow Matching
@@
In the fine stage, the model further refines this approximation by recovering fine-grained details $x^{i,f}$, represented as the residual between the original frame $x^{i}$ and the upsampled coarse component $\text{{Upsample}}(x^{i,c})$. A secondary flow-matching model predicts the vector field $v_{t}^{f}(x^{i,f},z^{i},x^{i,c};\theta_{\text{FM}}^{f})$, governing this process by leveraging both the features $z^{i}$ and the coarse component (with ground-truth coarse features $x^{i,c}$ during training and predicted values $\tilde{x}^{i,c}$ during inference) as conditional inputs. This hierarchical conditioning allows the fine model to focus on local details while preserving global coherence from the coarse stage.
@@
$\mathcal{L}_{\text{C2F-FM}}\!=\!\mathbb{E}_{t,x_{1}^{i,c},x^{i,c}}\big{\|}u_{t}^{c}-v_{t}^{c}(x^{i,c},z^{i};\theta_{\text{FM}}^{c})\big{\|}^{2}$
(3) $+\mathbb{E}_{t,x_{1}^{i,f},x^{i,f}}\big{\|}u_{t}^{f}\!-\!v_{t}^{f}(x^{i,f},z^{i},x^{i,c};\theta_{\text{FM}}^{f})\big{\|}^{2},$
@@
Classifier-free guidance (CFG) is a powerful technique to enhance the quality and controllability of generated outputs in flow matching and diffusion models (Ho and Salimans, 2022; Nichol and Dhariwal, 2021). In FELLE, we implement CFG through joint training of coarse and fine flow matching models using both conditional and unconditional objectives. During training, we randomly mask the speech prompt with probability $p_{\text{drop}}$ for unconditional learning, which enables each model to learn dual vector fields. At inference, guided vector fields are computed by linear blending:
(4) $\hat{v}_{t}^{\ast}(x^{\ast};\cdot)$ $=wv_{t}^{\ast}(x^{\ast},c;\theta_{\text{FM}}^{\ast})+(1-w)v_{t}^{\ast}(x^{\ast},\bar{c};\theta_{\text{FM}}^{\ast}),$
@@
4.3. Training Objective
In FELLE, we integrate the condition loss $\mathcal{L}_{\text{cond}}$ in addition to coarse-to-fine loss $\mathcal{L}_{\text{C2F-FM}}$. $\mathcal{L}_{\text{cond}}$ is a hybrid loss function that combines L1 and L2 norms, defined as $\mathcal{L}_{\text{cond}}=\|z_{i}-x_{i}\|_{1}+\|z_{i}-x_{i}\|_{2}^{2}$, for step $i$ to regularize the conditional input for flow matching. Additionally, we introduce a stop prediction module to the autoregressive language model. This module, during each step of generation, transforms the hidden state output by the language model into the probability of a stop signal through a linear layer and calculates the Binary Cross-Entropy loss $\mathcal{L}_{\text{stop}}$ for training. The model can automatically determine when to stop during the generation process without the need to preset length rules. The overall training objective is: $\mathcal{L}=\mathcal{L}_{\text{C2F-FM}}+\lambda\mathcal{L}_{\text{cond}}+\alpha\mathcal{L}_{\text{stop}},$ where $\lambda$ and $\alpha$ control the respective contributions of $\mathcal{L}_{\text{cond}}$ and $\mathcal{L}_{\text{stop}}$.
4.4. Inference
As illustrated in Figure 2(b), the inference process employs an autoregressive language model that progressively generates hidden representations based on textual and speaker prompts. At each step $i$, the computed latent state $z_{i}$ serves two key purposes. First, it provides conditional guidance for the coarse flow-matching module, facilitating the gradual transformation from the previous mel-spectrogram approximation $\tilde{x}^{i-1,c}$ to the current coarse structural estimate $\tilde{x}^{i,c}$. Following this coarse estimation phase, the integrated information of $\tilde{x}^{i,c}$ and $z_{i}$ drives the fine flow-matching module to produce the fined mel-spectrogram frame $\tilde{x}^{i,f}$. The final output frame $\tilde{x}^{i}$ emerges through the integration of these complementary coarse and refined predictions. Secondly, the latent state $z_{i}$ processed by the stop prediction module to compute the stop probability, which is compared against a predefined threshold to decide whether to terminate the process. The iterative generation continues until the stop criterion is satisfied, after which a neural vocoder converts the mel-spectrogram into the final speech waveform.
5. Experimental Setup
5.1. Dataset